OpenStudy (ragingsquirrel):

R - W R= 5X84+9 W= 2X5+30

OpenStudy (diyadiya):

\[5 \times 84 + 9= 429\]\[2 \times 5+ 3= 13\]\[R-W=429-13=416\] \( :) \)